OpenStudy (anonymous):

Explain this question fully and tell me what going on please! A boutique store owner purchases floor rugs for $45 each and marks them for sale so that he can sell then at a discount of 70% on his marked price and still make a profit of 80% on the selling price. (a) Calculate the marked price on these rugs. (b) Calculate the selling price of these rugs. (c) What profit is made on each of the rugs?

OpenStudy (anonymous):

what's*

OpenStudy (anonymous):

You start with the purchased price : $45. He wants 80% profit, ie, 45x80/100 = $36 as profit. So he wants to sell it at $45 + $36 = $81, which will become his 'sale price'. Now he wants to pretend he's giving a 70% discount when selling at $81. Which is another way of saying he wants to sell at 30% of his pretend price. So if his pretend price (called Marked Price or List Price) is P, then we have P x 30/100 = 81. Or, P = 81/ 0.3 = $270.

OpenStudy (anonymous):

I don't get the second paragraph @knowall

OpenStudy (anonymous):

Well, the problem states that he wants to 'offer' a discount of 70%. If he does not want to suffer a loss, this means he has to state a higher price (Marked Price), and then attract customers with a claim of "70% discount". Obviously that marked price has to be very high if he wants to sell at $81. The 2nd para describes how you get to that Marked price that he can list, and then offer 70% discount and still get a 80% profit.